# From Seeing to Doing: Bridging Reasoning and Decision for Robotic Manipulation

### v1

@@ 3.4 Self-Consistent Alignment for Spatial Understanding and Generation
High-quality SFT datasets enable VLMs to generate visual aids [66], yet these models struggle to understand the physical meaning of such annotations since coordinate spaces never appeared in pretraining data. The alignment between image coordinates and actual spatial positions presents a significant challenge. Therefore, we propose a self-consistency mechanism to further align FSD capabilities in spatial understanding and generation. We frame generation tasks inversely as understanding problems: if the forward task requires inferring visual trace $\boldsymbol{\tau}$ from an image $X_{v}$ and task instruction $X_{q}$, i.e. $(X_{v},X_{q})\rightarrow\boldsymbol{\tau}$, we construct the inverse task of predicting possible instructions given an image and visual traces $(X_{v},\boldsymbol{\tau})\rightarrow X_{q}$. This bidirectional approach helps the model comprehend spatial coordinates’ meanings and aligns coordinate space with image-text modalities, unifying visual aids as both understanding and generation signals while enhancing FSD spatial reasoning capabilities.
@@ 4 Training and Action Execution of FSD
Training: We follow the instruction tuning pipeline proposed by LLaVA-1.5 [33, 34]. As shown in Fig. 1, FSD’s architecture comprises an image encoder (CLIP-ViT-L-336px [11]), a linear projector, a language tokenizer, and a LLM (Vicuna-13B [71]). The image encoder processes images into tokens. These visual tokens are then projected into the same embedding space as language tokens through a two-layer linear projector. Only the projector and LLM weights are updated during fine-tuning while the vision encoder and tokenizer remain frozen. We built upon ASMv2 [55] as our foundation, which already incorporates basic relation conversation and reference grounding capabilities. The training process of FSD is divided into two stages: General Spatial Reasoning Enhancement: Using data from levels 1-3, we focus on improving the model’s embodied spatial reasoning capabilities. Following Yuan et al. [66] and Brohan et al. [3], we discovered that an appropriate data mixture is crucial for downstream performance. Joint training with mixed robotic and internet data ensures the model retains knowledge acquired during pre-training. Consequently, our instruction tuning utilizes a diverse 1.4M sample mixture including general visual question answering (VQA) data. This comprehensive training ensures FSD maintains robust general spatial knowledge while developing embodied capabilities. Visual Aids Generation and Understanding: Using data from levels 4-5 with the self-consistency mechanism, we specifically train visual aids generation and understanding abilities. FSD predicts a fixed set of 8 points for simplification when generating spatial visual traces. Additional training details and the summary of mixture datasets are provided in Appendix B.
@@
Action Execution: FSD can reason from initial or intermediate task steps, freely selecting needed visual aids. When using bounding boxes, we sample the center as the target point; with affordance points, we directly sample one point. For visual trace execution, we first generate 2D visual traces $\boldsymbol{\tau}$ and obtain preliminary depth information from depth cameras. Following the pinhole camera model, we employ depth-based back-projection to map these to 3D space, yielding $\boldsymbol{\tau}^{3d}=\{\mathbf{x}_{t}^{3d}\mid t=1,2,...,T\}$. Next, based on the spatial position of the first point $\mathbf{x}_{1}$, we query GraspNet’s [9] grasp candidates $G$ to match the nearest grasp pose $G^{*}$. For relatively fixed scenes, we may also use predetermined grasp poses. Subsequently, we optimize the path trajectory using gradient descent-based interpolation, generating complete motion trajectories in SE(3) space, enabling the robotic arm to follow the 3D visual trajectory. When using only spatial affordance, we utilize CuRobo [49] as the motion planner to determine execution trajectories $T$ based on the target position $\mathcal{P}$. More details are provided in Appendix C. Unlike methods such as LLARVA [39] and EmbodiedCOT [67] which also utilize visual auxiliary aids, FSD transforms prediction tasks into reasoning tasks, better leveraging visual-spatial common knowledge without requiring scenario-specific fine-tuning.
@@ 5 Visual Aids Generation Benchmark
Few datasets exist for evaluating visual aid generation, Where2Place [66] provides 100 real-world images from homes and offices but is limited to direct and simple language instructions. And no benchmarks for trajectory prediction. To address this gap, we propose the Visual Aids Generation Benchmark (VABench). we manually annotated 300 problems from real-world and simulation datasets (OXE [40], BridgeData [53], Droid [20], Libero [31]), requiring models to infer visual aids given only natural language instructions similar to everyday human commands. For spatial affordance (VABench-Point), we measure the proportion of points falling within target regions. For models that only output bounding boxes, we sample uniformly within boxes. For visual trace (VABench-VisualTrace), we compute MAE and RMSE between ground truth $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ and predictions $\hat{\boldsymbol{\tau}}=\{\hat{\mathbf{x}}_{t}\mid t=1,2,...,\hat{T}\}$: $\text{MAE}=\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|$, $\text{RMSE}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}}$. We interpolate when trajectory lengths differ and normalize coordinates to a 1000×1000 space for consistent evaluation. Since multiple valid solutions exist for each instruction, we simulated human evaluation standards by establishing detailed assessment criteria and added metrics with MLLM-based qualitative scoring (0-10) of visualized trajectories, named GPT Score. We provide detailed dataset information and evaluation procedures in Appendix D.
@@ 6 Experiments
We evaluated FSD across two dimensions: Seeing and Doing. For Seeing, we tested its general spatial reasoning and visual aids generation capabilities. For Doing, we conducted zero-shot manipulation experiments in both SimplerEnv [25] simulation and real-world xArm robotic platforms to assess its practical generalization performance.
@@ 6.1 Evaluation of Spatial Understanding and Reasoning Capabilities
Benchmarks and Baselines. General Spatial Reasoning Capabilities: We evaluated general spatial reasoning capabilities using five popular benchmarks: CVBench [51], BLINK [10], CRPE [55], SAT [44], and EmbSpatial-Bench [8]. These benchmarks encompass 15 subtasks measuring various spatial competencies including counting, spatial relationships, distance estimation and so on. We included two leading closed-source models: GPT-4o [15] and GPT-4V as performance reference. Subsequently, we conducted comparative analyses against other open-source spatial enhanced MLLMs, including LLaVA-1.5 [33], SAT-Dynamic [44], RoboPoint [66], and ASMv2 [55], all with 13B parameters. Object and Free Region Reference Capabilities: Following Yuan et al. [66], we assessed embodied spatial capabilities using the RoboRefIt [36] and Where2Place [66] benchmarks. We compared mainstream closed-source models and MLLM for enhancing spatial abilities (SpatialBot [4], SpaceLLaVA [5], and RoboBrain [16]). We used the proportion of predicted points within specified objects/regions as the accuracy metric. For models without point output support, we asked models to output bounding boxes of target objects/free space regions, then sampled evenly within these bounding boxes. Spatial Affordance and Visual Trace Capabilities: We utilized our VABench to evaluate the capabilities. We found very few models with this capability for Visual Trace prediction, so we trained an end-to-end prediction baseline model using a pre-trained DINOv2 [41] encoder coupled with a multi-layer transformer [52] architecture to predict visual trajectories, trained on the same visual trajectory data, named DINOv2 Predictor. We conducted this comparison to demonstrate the advantages of our reasoning-based FSD approach. More detailed descriptions of benchmarks and baselines are provided in Appendix E.
@@
CVBench CRPE SAT BLINK EmbSpatial Rank
@@
DINOv2 Predictor 128.32 117.49 2.17
FSD 78.26 63.44 5.49
   w/o SrCoT 99.53 80.06 4.00
   w/o Alignment 80.48 66.80 5.45
Table 4: SimplerEnv Evaluation on WidowX Robot The results of baselines are derived from [43]. ZS: zero-shot, FT: fine-tuning using BridgeData. Each task is tested 24 episodes.
Model Spoon→Towel Carrot→Plate Green→Yellow Eggplant→Basket Avg.
@@
RT-1-X [40] 16.7% 0% 20.8% 4.2% 8.3% 0% 0.0% 0% 1.1%
Octo-S [50] 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 87.5% 56.9% 30.0%
OpenVLA [21] 4.1% 0% 33.3% 0% 12.5% 0% 8.3% 4.1% 1.0%
RoboVLM (ZS) [24] 37.5% 20.8% 33.3% 25.0% 8.3% 8.3% 0.0% 0% 13.5%
RoboVLM (FT) [24] 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 58.3% 58.3% 31.3%
SpatialVLA (ZS) [43] 25.0% 20.8% 41.7% 20.8% 58.3% 25.0% 79.2% 70.8% 34.4%
SpatialVLA (FT) [43] 20.8% 16.7% 29.2% 25.0% 62.5% 29.2% 100% 100% 42.7%
FSD 58.3% 41.6% 58.3% 50.0% 91.6% 33.3% 37.5% 37.5% 54.1%
@@
FSD demonstrates breakthrough performance in visual aid generation. As shown in Table Table 3(b), FSD significantly outperforms all baseline methods in generating precise spatial affordances and visual traces. For VABench-Point, FSD achieves 61.82% accuracy—three times higher than RoboPoint (19.09%), while for VABench-VisualTrace, FSD demonstrates significantly lower error rates and better GPT Score compared to the DINOv2 Predictor. Ablation studies confirm the critical contributions of both SrCoT and self-consistency alignment, validating our reasoning-based approach where step-by-step spatial analysis enables more accurate predictions than purely data-driven methods. We provide detailed comparisons with RoboBrain [16] in Appendix I.
@@
Visualization of visual aids generated by FSD We present examples of generation in Fig. 5. FSD is able to adapt to diverse scenes, perspectives, and tasks, and the three types of visual traces generated can well match the task instructions. We also presented the reasoning process of FSD in Fig. 6, where the CoT provides clear guidance for the results. More visualizations are in Appendix F.
6.2 Evaluate the Decision Capability
@@
SimplerEnv Simulation SimplerEnv [25] is a simulation environment specifically designed for evaluating real-world robotic manipulation. It provides a standardized platform, emphasizing reproducibility with real-world scenarios. We utilized FSD to generate visual traces and perform zero-shot deployment on the WidowX robotic arm. We compared its performance against state-of-the-art general manipulation VLA, including Octo [50], OpenVLA [21], RoboVLM [24] and SpatialVLA [43]. The results are presented in Table 4. FSD surpasses the baseline in most tasks without FT with a 54.1% success rate, demonstrating the strong generalization ability of FSD. Notably, FSD achieves the best average performance, even outperforming VLA models that even outperforming specially fine-tuned VLA models. This superior generalization capability demonstrates that visual aids representations provide better zero-shot transfer compared to direct action prediction approaches, bridging the gap between perception and control more effectively.
@@
Real World Robot Evaluation As shown in Fig. 7, we conducted zero-shot tests with FSD on an xArm 6 robot across 8 tabletop manipulation tasks. The setup included an Intel RealSense L515 LiDAR camera. To test the capabilities of different visual aids, we used visual trace for the sponge and folding tasks, while affordance points were used for other tasks. We compared against RoboPoint and GPT-4o baselines, which were limited to predicting only start/end points rather than full trajectories. Robopoint often makes incorrect predictions in tasks involving spatial understanding. Under zero-shot conditions, FSD achieved 72% success rate, outperforming the strongest baseline by more than 30%. Notably, FSD successfully completed complex tasks with visual trace generation, e.g. cloth folding, which was beyond baseline capabilities. Full results are presented in Fig. 7 and we refer to Appendix G for detailed setup and more visualizations.

### v2

@@ 3.4 Self-Consistent Alignment for Spatial Understanding and Generation
High-quality SFT datasets enable VLMs to generate visual aids Yuan et al. (2024b), yet these models struggle to understand the physical meaning of such annotations since coordinate spaces never appeared in pretraining data. The alignment between image coordinates and actual spatial positions presents a significant challenge. Therefore, we propose a self-consistency mechanism to further align FSD capabilities in spatial understanding and generation. We frame generation tasks inversely as understanding problems: if the forward task requires inferring visual trace $\boldsymbol{\tau}$ from an image $X_{v}$ and task instruction $X_{q}$, i.e. $(X_{v},X_{q})\rightarrow\boldsymbol{\tau}$, we construct the inverse task of predicting possible instructions given an image and visual traces $(X_{v},\boldsymbol{\tau})\rightarrow X_{q}$. This bidirectional approach helps the model comprehend spatial coordinates’ meanings and aligns coordinate space with image-text modalities, unifying visual aids as both understanding and generation signals while enhancing FSD spatial reasoning capabilities.
@@ 4 Training and Action Execution of FSD
Training: We follow the instruction tuning pipeline proposed by LLaVA-1.5 Liu et al. (2023c, 2024b). As shown in Fig. 1, FSD’s architecture comprises an image encoder (CLIP-ViT-L-336px Gao et al. (2024)), a linear projector, a language tokenizer, and a LLM (Vicuna-13B Zheng et al. (2023b)). The image encoder processes images into tokens. These visual tokens are then projected into the same embedding space as language tokens through a two-layer linear projector. Only the projector and LLM weights are updated during fine-tuning while the vision encoder and tokenizer remain frozen. We built upon ASMv2 Wang et al. (2025) as our foundation, which already incorporates basic relation conversation and reference grounding capabilities. The training process of FSD is divided into two stages: General Spatial Reasoning Enhancement: Using data from levels 1-3, we focus on improving the model’s embodied spatial reasoning capabilities. Following Yuan et al. (2024b) and Brohan et al. (2023), we discovered that an appropriate data mixture is crucial for downstream performance. Joint training with mixed robotic and internet data ensures the model retains knowledge acquired during pre-training. Consequently, our instruction tuning utilizes a diverse 1.4M sample mixture including general visual question answering (VQA) data. This comprehensive training ensures FSD maintains robust general spatial knowledge while developing embodied capabilities. Visual Aids Generation and Understanding: Using data from levels 4-5 with the self-consistency mechanism, we specifically train visual aids generation and understanding abilities. FSD predicts a fixed set of 8 points for simplification when generating spatial visual traces. Additional training details and the summary of mixture datasets are provided in Appendix B.
@@
Action Execution: FSD can reason from initial or intermediate task steps, freely selecting needed visual aids. When using bounding boxes, we sample the center as the target point; with affordance points, we directly sample one point. For visual trace execution, we first generate 2D visual traces $\boldsymbol{\tau}$ and obtain preliminary depth information from depth cameras. Following the pinhole camera model, we employ depth-based back-projection to map these to 3D space, yielding $\boldsymbol{\tau}^{3d}=\{\mathbf{x}_{t}^{3d}\mid t=1,2,...,T\}$. Next, based on the spatial position of the first point $\mathbf{x}_{1}$, we query GraspNet’s Fang et al. (2020) grasp candidates $G$ to match the nearest grasp pose $G^{*}$. For relatively fixed scenes, we may also use predetermined grasp poses. Subsequently, we optimize the path trajectory using gradient descent-based interpolation, generating complete motion trajectories in SE(3) space, enabling the robotic arm to follow the 3D visual trajectory. When using only spatial affordance, we utilize CuRobo Sundaralingam et al. (2023) as the motion planner to determine execution trajectories $T$ based on the target position $\mathcal{P}$. More details are provided in Appendix C. Unlike methods such as LLARVA Niu et al. (2024) and EmbodiedCOT Zawalski et al. (2024) which also utilize visual auxiliary aids, FSD transforms prediction tasks into reasoning tasks, better leveraging visual-spatial common knowledge without requiring scenario-specific fine-tuning.
@@ 5 Visual Aids Generation Benchmark
Few datasets exist for evaluating visual aid generation, Where2Place Yuan et al. (2024b) provides 100 real-world images from homes and offices but is limited to direct and simple language instructions. And no benchmarks for trajectory prediction. To address this gap, we propose the Visual Aids Generation Benchmark (VABench). we manually annotated 300 problems from real-world and simulation datasets (OXE O’Neill et al. (2023), BridgeData Walke et al. (2023), Droid Khazatsky et al. (2024) and Libero Liu et al. (2023a)), requiring models to infer visual aids given only natural language instructions similar to everyday human commands. For spatial affordance (VABench-Point), we measure the proportion of points falling within target regions. For models that only output bounding boxes, we sample uniformly within boxes. For visual trace (VABench-VisualTrace), we compute MAE and RMSE between ground truth $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ and predictions $\hat{\boldsymbol{\tau}}=\{\hat{\mathbf{x}}_{t}\mid t=1,2,...,\hat{T}\}$: $\text{MAE}=\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|$, $\text{RMSE}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}}$. We interpolate when trajectory lengths differ and normalize coordinates to a 1000×1000 space for consistent evaluation. Since multiple valid solutions exist for each instruction, we simulated human evaluation standards by establishing detailed assessment criteria and added metrics with MLLM-based qualitative scoring (1-10) of visualized trajectories, named GPT Score. We provide detailed dataset information and evaluation procedures in Appendix D.
@@ 6 Experiments
We evaluated FSD across two dimensions: Seeing and Doing. For Seeing, we tested its general spatial reasoning and visual aids generation capabilities. For Doing, we conducted zero-shot manipulation experiments in both SimplerEnv Li et al. (2024e) simulation and real-world xArm robotic platforms to assess its practical generalization performance.
@@ 6.1 Evaluation of Spatial Understanding and Reasoning Capabilities
Benchmarks and Baselines. General Spatial Reasoning Capabilities: We evaluated general spatial reasoning capabilities using five popular benchmarks: CVBench Tong et al. (2024), BLINK Fu et al. (2024), CRPE Wang et al. (2025), SAT Ray et al. (2024), and EmbSpatial-Bench Du et al. (2024). These benchmarks encompass 15 subtasks measuring various spatial competencies including counting, spatial relationships, distance estimation and so on. We included two leading closed-source models: GPT-4o Hurst et al. (2024) and GPT-4V as performance reference. Subsequently, we conducted comparative analyses against other open-source spatial enhanced MLLMs, including LLaVA-1.5 Liu et al. (2023c), SAT-Dynamic Ray et al. (2024), RoboPoint Yuan et al. (2024b), and ASMv2 Wang et al. (2025), all with 13B parameters. Object and Free Region Reference Capabilities: Following Yuan et al. (2024b), we assessed embodied spatial capabilities using the RoboRefIt Lu et al. (2023) and Where2Place Yuan et al. (2024b) benchmarks. We compared mainstream closed-source models and MLLM for enhancing spatial abilities (SpatialBot Cai et al. (2024), SpaceLLaVA Chen et al. (2024), and RoboBrain Ji et al. (2025)). We used the proportion of predicted points within specified objects/regions as the accuracy metric. For models without point output support, we asked models to output bounding boxes of target objects/free space regions, then sampled evenly within these bounding boxes. Spatial Affordance and Visual Trace Capabilities: We utilized our VABench to evaluate the capabilities. We found very few models with this capability for Visual Trace prediction, so we trained an end-to-end prediction baseline model using a pre-trained DINOv2 Oquab et al. (2023) encoder coupled with a multi-layer transformer Vaswani et al. (2017) architecture to predict visual trajectories, trained on the same visual trajectory data, named DINOv2 Predictor. We conducted this comparison to demonstrate the advantages of our reasoning-based FSD approach. More detailed descriptions of benchmarks and baselines are provided in Appendix E.
@@
Figure 6: Visualization of FSD executing tasks. The first six rows are real-world experiments, and the last four rows are from SIMPLEREnv.
Figure 7: An example and visualization of FSD for generating visual trace.
Figure 8: An example and visualization of FSD for generating affordance box and points.
@@
CVBench CRPE SAT BLINK EmbSp Rank
@@
GPT4o 136.13 113.53 4.37
DINOv2 Predictor 128.32 117.49 4.01
FSD 78.26 63.44 6.21
   w/o SrCoT 99.53 80.06 5.07
   w/o Alignment 80.48 66.80 5.92
Table 4: SimplerEnv Evaluation on WidowX Robot. The results of baselines are derived from Qu et al. (2025). ZS: zero-shot, FT: fine-tuning using BridgeData. Each task is tested 24 episodes.
[…]
@@
RT-1-X O’Neill et al. (2023) 16.7% 0% 20.8% 4.2% 8.3% 0% 0.0% 0% 1.1%
Octo-S Team et al. (2024) 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 87.5% 56.9% 30.0%
OpenVLA Kim et al. (2024) 4.1% 0% 33.3% 0% 12.5% 0% 8.3% 4.1% 1.0%
RoboVLM (ZS) Li et al. (2024d) 37.5% 20.8% 33.3% 25.0% 8.3% 8.3% 0.0% 0% 13.5%
RoboVLM (FT) Li et al. (2024d) 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 58.3% 58.3% 31.3%
SpatialVLA (ZS) Qu et al. (2025) 25.0% 20.8% 41.7% 20.8% 58.3% 25.0% 79.2% 70.8% 34.4%
SpatialVLA (FT) Qu et al. (2025) 20.8% 16.7% 29.2% 25.0% 62.5% 29.2% 100% 100% 42.7%
FSD 58.3% 41.7% 58.3% 50.0% 91.7% 33.3% 37.5% 37.5% 40.6%
@@
FSD demonstrates breakthrough performance in visual aid generation. As shown in Table Table 3(b), FSD significantly outperforms all baseline methods in generating precise spatial affordances and visual traces. For VABench-Point, FSD achieves 61.82% accuracy—three times higher than RoboPoint (19.09%), while for VABench-VisualTrace, FSD demonstrates significantly lower error rates and better GPT Score compared to the DINOv2 Predictor. Ablation studies confirm the critical contributions of both SrCoT and self-consistency alignment, validating our reasoning-based approach where step-by-step spatial analysis enables more accurate predictions than purely data-driven methods. We provide detailed comparisons with RoboBrain Ji et al. (2025) in Appendix I.
@@
Visualization of visual aids generated by FSD We present examples of generation in Fig. 5. FSD is able to adapt to diverse scenes, perspectives, and tasks, and the three types of visual traces generated can well match the task instructions. We also presented the reasoning process of FSD in Fig. 9, where the CoT provides clear guidance for the results. Subsequently, we selected several specific examples to showcase the complete outputs of FSD in Fig. 8 and Fig. 7. More visualizations are in Appendix F.
6.2 Evaluation of the Decision Capability
@@
SimplerEnv Simulation SimplerEnv Li et al. (2024e) is a simulation environment specifically designed for evaluating real-world robotic manipulation. It provides a standardized platform, emphasizing reproducibility with real-world scenarios. We utilized FSD to generate visual traces and perform zero-shot deployment on the WidowX robotic arm. We compared its performance against state-of-the-art general manipulation VLA, including Octo Team et al. (2024), OpenVLA Kim et al. (2024), RoboVLM Li et al. (2024d) and SpatialVLA Qu et al. (2025). The results are presented in Table 4. FSD surpasses the baseline in most tasks without FT with a 40.6% success rate, demonstrating the strong generalization ability of FSD. Notably, FSD achieves the best zero-shot average performance and is competitive with specially fine-tuned VLA models. This superior generalization capability demonstrates that visual aids representations provide better zero-shot transfer compared to direct action prediction approaches, bridging the gap between perception and control more effectively.
@@
Real World Robot Evaluation As shown in Fig. 10, we conducted zero-shot tests with FSD on an xArm 6 robot across 8 tabletop manipulation tasks. The setup included an Intel RealSense L515 LiDAR camera. To test the capabilities of different visual aids, we used visual trace for the sponge and folding tasks, while affordance points were used for other tasks. We compared against RoboPoint and GPT-4o baselines, which were limited to predicting only start/end points rather than full trajectories. Robopoint often makes incorrect predictions in tasks involving spatial understanding. Under zero-shot conditions, FSD achieved 72% success rate, outperforming the strongest baseline by more than 30%. Notably, FSD successfully completed complex tasks with visual trace generation, e.g. cloth folding, which was beyond baseline capabilities. Full results are presented in Fig. 10 and visualization of control tasks executed using FSD in Fig. 6. We refer to Appendix G for detailed setup.